# Exploring the Impact of HAPS-RIS on UAV-Based Networks: a Novel Network Architecture

### v2

@@ Abstract
In this paper, we propose a network architecture where two types of aerial infrastructures together with a ground station provide connectivity to a remote area. A high altitude platform station (HAPS) is equipped with reconfigurable intelligent surface (RIS), so-called HAPS-RIS, to be exploited to assist the unmanned aerial vehicle (UAV)-based wireless networks. A key challenge in such networks is the restricted number of UAVs, which limits full coverage and leaves some users unsupported. To tackle this issue, we propose a hierarchical bilevel optimization framework including a leader and a follower problem. The users served by HAPS-RIS are in a zone called HAPS-RIS zone and the users served by the UAVs are in another zone called UAV zone. In the leader problem, the goal is to establish the zone boundary and practical RIS phase shift design that maximizes the number of users covered by HAPS-RIS while ensuring that users in this zone meet their rate requirements. This is achieved through our proposed relaxation method and algorithm based on the RIS clustering concept. The follower problem focuses on minimizing the number of UAVs required, ensuring that the rate requirements of the users in the UAV zone are met. This is addressed through our proposed dynamic method adopting k-means clustering technique with adaptive UAV selection. Our study reveals that increasing the number of RIS elements significantly decreases the number of required UAVs.
Index Terms:
RIS, NTNs, HAPS, UAV, Network architecture.
@@ I Introduction
A primary aim of future wireless networks is to ensure universal connectivity, offering reliable access to everyone and everything, anywhere and anytime, at an affordable price. This includes extending coverage to remote areas and the areas where establishing terrestrial infrastructure is not feasible or affordable [1]. To achieve this aim, unmanned aerial vehicles (UAVs), can play a vital role, operating as aerial base stations in regions where there is no cellular infrastructures or the areas where building a cellular infrastructure is very expensive. This includes scenarios like remote regions and disaster-stricken areas where ground communications may be destroyed, or any area where it is not affordable for network operators to establish terrestrial infrastructures [3, 2, 4]. However, one of the main challenges in the UAV-assisted networks is the limitation on the number of UAVs that can be deployed, which leads to limitation on the coverage. Deploying a network of multiple UAVs to provide full coverage over a specific area requires substantial investment from network providers. Therefore, it is essential to minimize the number of UAVs to improve the practical deployment of multi-UAV networks [5]. With the limitation on the number of UAVs, full coverage can not be achieved and portion of users are left unsupported. One solution to tackle this issue is to use other types of non-terrestrial infrastructures as transmitters or relays. However, this approach is neither low-cost nor low-complexity, as it requires the implementation of multiple antennas on the non-terrestrial platform [6]. Another option is to utilize available terrestrial infrastructures despite their far distance to support the UAV-based network. This can be done by connecting them to a non-terrestrial platform equipped with a reconfigurable intelligent surface (RIS). This system is called non-terrestrial (NT)-RIS, a smart reflective layer acting as an intelligent intermediary for signal reflection. Exploiting NT-RIS has lower cost and less complexity than the aforementioned approach, using other types of non-terrestrial infrastructures [6]. A considerable amount of research has been dedicated to exploring the advantages of implementing NT-RIS in wireless networks, in recent years, see [6, 7, 8, 9] and the related references therein. High altitude platform station (HAPS)-RIS stands out as a promising option for NT-RIS deployment, offering advantages over other non-terrestrial platforms like satellite-RIS and UAV-RIS [8, 9]. Accordingly, the main research question that arises is “Can HAPS-RIS effectively address the mentioned challenge which is the limitation on the number of UAVs?
@@
In [9], the link budget of the RIS-enabled non-terrestrial platforms are studied and location of the HAPS-RIS is optimized. This study finds that HAPS, with its large area for RIS deployment, outperforms other aerial platforms like UAVs and low earth orbit (LEO) satellites in terms of signal strength. In [10], the authors present a scenario where HAPS-RIS is exploited to assist the network supported by terrestrial base stations and improve its service in urban areas. The users which are not covered with the terrestrial base stations, can be connected to the dedicated control station (CS) through the HAPS-RIS system. The paper introduces optimization problems including throughput maximization and minimizing the number of RIS elements by designing the RIS phase shifts and the power allocation of the CS. In [11], the authors propose resource efficient approach, for the same scenario as [10], that aims to connect the maximum number of users while also minimizing the overall power usage of the CS and the RIS. In [12], the authors present a method for channel estimation in full-duplex backhaul communication from base station to core netwrok supported by HAPS-RIS. Although the focus is on channel estimation, this work is related to our research as it studies the performance of a network supported by HAPS-RIS. In [13], a multi-objective optimization problem is proposed where two unconnected infrastructures are connected though an aerodynamic HAPS-RIS. A closed-form solution for the RIS phase shifts are obtained to maximize the cascade channel gain, eliminate Doppler spread, and minimize the upper limit of delay spread.
@@
While there has been considerable progress in exploiting HAPS-RIS to improve wireless networks, a critical gap remains in applying this technology to UAV-assisted networks. To the best of our knowledge there is no work which study the integration of HAPS-RIS technology and UAV-assisted networks where they serve as aerial base stations. In summary, this paper addresses the afromentioend gap in the literature by highlighting the following contributions.
@@
  • We introduce a novel network architecture where a supportive network, exploiting HAPS-RIS, is established to solve one of the major challenges in the UAV-based network which is limitation on the number of needed UAVs.
@@
  • We mathematically model the heterogeneous zones, i.e., UAV zone and HAPS-RIS zone for user clustering and associate the users of the zones to their allocated infrastructures, i.e., UAV or HAPS-RIS.
@@
  • We propose a novel hierarchical bilevel optimization problem, including a leader and a follower problem. In the leader one, we obtain the zone boundary and a practical design for the RIS phase shifts to maximize the number of users supported by HAPS-RIS while satisfying the rate requirements for each user in the HAPS-RIS zone. We relax and redesign the leader problem into an optimization problem with the lower computational complexity. In the follower optimization problem, we aim to minimize the number of UAVs while meeting the rate requirement for each user in the UAV zone. To solve the follower problem, we propose a dynamic method adopting $k$-means clustering technique with adaptive UAV selection. Our findings demonstrate that enhancing the number of RIS elements can significantly reduce the required number of UAVs, revealing a key strategy for optimizing aerial operations.
II System Model
@@
We consider a downlink transmission scenario, shown in Fig. 1, where heterogeneous non-terrestrial infrastructures are exploited to bring connectivity to a remote area. The set of stationary users $\mathcal{I}$ is defined as $\mathcal{I}$={1, …, $I$}, where $i\in\mathcal{I}$ represents an individual user indexed by $i$. The users are uniformly distributed in a circular area where there is no nearby terrestrial infrastructure. The set of UAVs $\mathcal{J}$ is defined as $\mathcal{J}$ = {1, …, $J$}, where $j\in\mathcal{J}$ represents an individual UAV indexed by $j$. In this scenario, the UAVs remain stationary, hovering in place to provide service to the fixed users. The aim is minimizing the number of required UAVs with exploiting terrestrial CS and HAPS-RIS while providing full coverage with meeting the minimum rate requirements. In HAPS-RIS scenarios, CS uses high gain directional antenna, creating a dominated line of sight (LoS) connection with HAPS [8, 9, 10, 11, 12, 13]. We consider UAVs and CS operate independently to provide connectivity and they are not connected to each other. All UAVs can be connected to the core network via a backhaul link. Due to heavy blockage, there is no direct link between the users and the CS; communication instead occurs through a cascaded channel through HAPS-RIS. The set of RIS elements $\mathcal{M}$ is defined as $\mathcal{M}$ = {1, …, $M$}, where $m\in\mathcal{M}$ represents the RIS element indexed by $m$. Users can be connected to the network either through the UAV system or the HAPS-RIS setup, but not through both at the same time. The coverage area is a circle with the radius $R_{0}$ with the centre of $(x_{0},y_{0},z_{0})$, the CS location is $(x^{\rm{CS}},y^{\rm{CS}},z^{\rm{CS}})$, the location of the HAPS is $(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}})$, and the location of the UAV $j$ is $(x_{j}^{\rm{UAV}},y_{j}^{\rm{UAV}},z_{j}^{\rm{UAV}})$ in the Cartesian coordinate system. As can be seen in Fig. 1, there are two zones in the coverage area, i.e., the inner circle with the radius $R$ and the ring around it. The UAV zone is the area where the users are covered by UAVs and the HAPS-RIS zone is the ring area between the inner and outer circles, in which the users are covered by HAPS-RIS. The border which seperates the UAV zone and the HAPS-RIS zone is called zone boundary. $\mathcal{B}$ is the set which includes the users in the UAV zone. Thus, $\mathcal{C}$$\mathcal{I-B}$ is the set which includes the users in the HAPS-RIS zone. The set of equally spaced orthogonal subcarriers $\mathscr{L}$ is defined as $\mathscr{L}$={1, …, $L^{\rm{tot}}$}, where $l\in\mathscr{L}$ represents the subcarrier indexed by $l$. Let us consider the available bandwidth of $BW$ which is divided into $L^{\rm tot}$. Therefore, the bandwidth allocated to each orthogonal subcarrier is $B_{l}=\frac{BW}{L^{\rm{tot}}}$.
Figure 1: The Proposed System Model
III Problem Formulation and Proposed Solution
@@
III-A HAPS-RIS Zone
@@
$h^{\rm{CS}}_{im}=\sqrt{G^{\rm{CS}}G_{i}\left(\mathcal{L}^{\rm{CS-RIS}}_{m}\right)^{-1}\left(\mathcal{L}^{\rm{RIS-user}}_{im}\right)^{-1}},\forall m,\forall i\in\mathcal{C},$ (1)
where $G^{\rm{CS}}$ is the antenna gain of the CS and $G_{i}$ is the antenna gain of the user $i$. $(x^{\rm{CS}},y^{\rm{CS}},z^{\rm{CS}})$, $\left(x_{i},y_{i},z_{i}\right)$, $(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}})$, $\left(x^{\rm{RIS}}_{m},y^{\rm{RIS}}_{m},z^{\rm{RIS}}_{m}\right)$ are the CS, user $i$, HAPS, and RIS element $m$, coordinates, respectively. Using the Friis model, [14], the path loss between CS and RIS element $m$ can be formulated as
$\begin{array}[]{l}\mathcal{L}^{\rm{CS},\rm{RIS}}_{m}=(\frac{4\pi f_{\rm c}}{c})^{2}\times\\ \left(\left(x^{\rm{RIS}}_{m}-x^{\rm{CS}})^{2}+(y^{\rm{RIS}}_{m}-y^{\rm{CS}})^{2}+(z^{\rm{RIS}}_{m}-z^{\rm{CS}}\right)^{2}\right),\ \forall m,\end{array}$ (2)
and the path loss between RIS element $m$ and the user $i$ can be formulated as
$\begin{array}[]{l}\mathcal{L}^{\rm{RIS},\rm{user}}_{im}=\left(\frac{4\pi f_{\rm{c}}}{c}\right)^{2}\times\\ \left((x^{\rm{RIS}}_{m}-x_{i})^{2}+(y^{\rm{RIS}}_{m}-y_{i})^{2}+(z^{\rm{RIS}}_{m}-z_{i})^{2}\right),\forall m,\forall i\in\mathcal{C},\end{array}$ (3)
@@
where $L^{\rm{CS}}$=$\frac{L^{\rm{tot}}}{2}$ is available subcarriers, $\gamma^{l,\rm{CS}}_{i}$, the signal to noise ratio (SNR) of the received signal at user $i$ and subcarrier $l$, can be obtained as
@@
where $M$ is the total number of RIS elements, $N_{0}$ is the noise power spectral density, and $P^{l,\rm{CS}}$ is the CS transmit power allocated uniformly to the subcarrier $l$. $\rho^{l,\rm{CS}}_{im}$ is the binary variable which demonstrates the user $i$ is assigned to RIS element $m$ over subcarrier $l$. Considering $\vartheta_{im}$ the reflection coefficient of the RIS element $m$ corresponding to the user $i$ as
@@
Definition 1: Similar to the approach in [11], where the justification is elaborated in detail, we consider each user in the HAPS-RIS zone is allocated to its own group of subcarriers and RIS elements. The group of RIS elements which are allocated to a user is called RIS cluster. Given this assumption, $\rho^{l,\rm{CS}}_{im}$ can be obtained based on the value of $R$, the radius of the inner circle, which determines the zone boundary, elaborated in Algorithm 1. Taking all these factors into account, the leader optimization problem can be formulated as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{OP 1}:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \max_{{R},\vartheta_{im}}\ U^{\rm{HAPS}}}$ (7a)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm{s}}.{\rm{t}}.:}\leavevmode\nobreak\ \leavevmode\nobreak\ r^{\rm{CS}}_{i}(R,\vartheta_{im})\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},}$ (7b)
@@
[…]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm{s}}.{\rm{t}}.:}\leavevmode\nobreak\ \leavevmode\nobreak\ r^{\rm{CS}}_{i}(R)\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},}$ (8b)
where $\phi^{\ast}_{m}=\xi_{0}+\omega_{0}+2k\pi,\leavevmode\nobreak\ \forall m$ and $k$ can be an integer value.
@@
In the proposed geometrical case, the dimensions of RIS can be negligible compared to the altitude of HAPS, as it does not have much effect on the performance. Hence, without loss of generality, we relax the location of the RIS elements as $\begin{matrix}\left(x^{\rm{RIS}}_{m},y^{\rm{RIS}}_{m},z^{\rm{RIS}}_{m}\right)\approx(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}}),&\forall m\end{matrix}.$  [^1] The dimensions of the RIS are in the range of 100 meters while the distance between the HAPS-RIS and the terrestrial network is in the range of more than 20 kilometers. The dimensions of the RIS and the coverage area is much less than the range of the distance between HAPS-RIS and the terrestrial components. Accordingly, without loss of generality, we can relax the corresponding phases to fixed values as $\xi^{\rm{CS}}_{m}\approx\xi_{0},\forall m$ and $\omega_{im}\approx\omega_{0},\forall i\in\mathcal{C},\forall m,$ so the reflection gain of RIS element $m$ can be relaxed as $\vartheta_{im}\approx{\tilde{\vartheta}_{m}},\forall i\in\mathcal{C},\forall m$ where
@@
As (9) is independent from user index, $r^{\rm{CS}}_{i}(R,\tilde{\vartheta}_{m})$ can be relaxed into its upper bound by considering $\phi^{\ast}_{m}=$$\arg\max_{\phi_{m}}$${\tilde{\vartheta}_{m}}=\xi_{0}+\omega_{0}+2k\pi,\forall m$. Accordingly, (7b) is relaxed into (8b) and hence OP 1 is relaxed into OP 1-a. The proposed technique significantly reduces the computational complexity by eliminating the need to individually optimize the phase shifts for each RIS element. Given a massive number of RIS elements in HAPS-RIS scenarios, ranging from $10^{5}$ to $10^{8}$, see Section. IV, the proposed design can pave the way for scalability and hence provides practical implementation.
@@
[…]
${{\rm{s}}.{\rm{t}}.:}\leavevmode\nobreak\ \leavevmode\nobreak\ r^{\rm{CS}}_{i}(R)\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},$ (10b)

### v3

@@ Abstract
In this paper, we propose a novel network architecture where two types of aerial infrastructures together with a ground station provide connectivity to a remote area. A high altitude platform station (HAPS) is equipped with reconfigurable intelligent surface (RIS), called HAPS-RIS, to be exploited to assist the unmanned aerial vehicle (UAV)-based wireless networks. A key challenge in such networks is the restricted number of UAVs, which limits full coverage and leaves some users unsupported. To tackle this issue, we propose a hierarchical bilevel optimization framework including a leader and a follower problem. The users served by HAPS-RIS are in a zone called the HAPS-RIS zone and the users served by the UAVs are in another zone called the UAV zone. In the leader problem, the goal is to establish the zone boundary and practical RIS phase shift design that maximizes the number of users covered by HAPS-RIS while ensuring that users in this zone meet their rate requirements. This is achieved through our proposed practical relaxation method and the proposed dynamic radius-based zone association with RIS clustering (DyRaZARC) technique. The follower problem focuses on minimizing the number of UAVs required, ensuring that the rate requirements of the users in the UAV zone are met. This is addressed through our proposed novel geometry-informed machine learning method, called k-means adaptive dynamic UAV selection (KADUS) technique. Our study reveals that increasing the number of RIS elements significantly decreases the number of required UAVs.
@@ I Introduction
Aprimary aim of future wireless networks is to ensure universal connectivity, offering reliable access to everyone and everything, anywhere and anytime, at an affordable price. This includes extending coverage to remote areas and the areas where establishing terrestrial infrastructure is not feasible or affordable [1]. To achieve this aim, unmanned aerial vehicles (UAVs), can play a vital role, operating as aerial base stations in regions where there is no cellular infrastructure or the areas where building a cellular infrastructure is very expensive. This includes scenarios like remote regions and disaster-stricken areas where ground communications may be destroyed, or any area where it is not affordable for network operators to establish terrestrial infrastructure [3, 2, 4]. However, one of the main challenges in the UAV-assisted networks is the limitation on the number of UAVs that can be deployed, which leads to limitation on the coverage. Deploying a network of multiple UAVs to provide full coverage over a specific area requires significant investment by network providers. Therefore, it is essential to minimize the number of UAVs to improve the practical deployment of multi-UAV networks [5]. With the limitation on the number of UAVs, full coverage cannot be achieved and a portion of users are left unsupported. One solution to tackle this issue is to use other types of non-terrestrial infrastructures as transmitters or relays. However, this approach is neither low-cost nor low-complexity, as it requires the implementation of multiple antennas on the non-terrestrial networks (NTN) [6]. Another option is to utilize available terrestrial infrastructures despite their far distance to support the UAV-based network. This can be done by connecting them to NTN equipped with a reconfigurable intelligent surface (RIS). This system is called non-terrestrial (NT)-RIS, a smart reflective layer acting as an intelligent intermediary for signal reflection. Exploiting NT-RIS has lower cost and less complexity than the aforementioned approach, using other types of non-terrestrial infrastructures [6]. A considerable amount of research has been dedicated to exploring the advantages of implementing NT-RIS in wireless networks, in recent years, see [6, 7, 8, 9] and the related references therein. High altitude platform station (HAPS)-RIS stands out as a promising option for NT-RIS deployment, offering advantages over other NT-RIS like satellite-RIS and UAV-RIS [8, 9]. Accordingly, the main research question that arises is “Can HAPS-RIS effectively address the mentioned challenge which is the limitation on the number of UAVs?
@@
In [9], the link budget of the RIS-enabled NTN are studied and location of the HAPS-RIS is optimized. This study finds that HAPS, with its large area for RIS deployment, outperforms other aerial platforms like UAVs and low earth orbit (LEO) satellites in terms of signal strength. In [10], the authors present a scenario where HAPS-RIS is exploited to assist the network supported by terrestrial base stations and improve its service in urban areas. The users which are not covered with the terrestrial base stations, can be connected to the dedicated control station (CS) through the HAPS-RIS system. The paper introduces optimization problems including throughput maximization and minimizing the number of RIS elements by designing the RIS phase shifts and the power allocation of the CS. In [11], the authors propose resource efficient approach, for the same scenario as [10], that aims to connect the maximum number of users while also minimizing the overall power usage of the CS and the RIS. In [12], a multi-objective optimization problem is proposed where two unconnected infrastructures are connected though an aerodynamic HAPS-RIS. A closed-form solution for the RIS phase shifts are obtained to maximize the cascade channel gain, eliminate Doppler spread, and minimize the upper limit of delay spread.
@@
While there has been considerable progress in exploiting HAPS-RIS to improve wireless networks, a critical gap remains in applying this technology to UAV-assisted networks. To the best of our knowledge there is no work which studies the integration of HAPS-RIS technology and UAV-assisted networks where they serve as aerial base stations. In summary, this paper addresses the aforementioned gap in the literature by highlighting the following contributions.
@@
  • We introduce a novel network architecture where a supportive network exploiting HAPS-RIS is established to address a key challenge in UAV-based networks, reducing the number of required UAVs.
@@
  • We mathematically model the heterogeneous zones, i.e., UAV zone and HAPS-RIS zone for user clustering and associate zone users with their allocated infrastructure, i.e., UAV or HAPS-RIS.
@@
  • We propose a novel hierarchical bilevel optimization problem, including a leader and a follower problem. In the leader one, we obtain the zone boundary and a practical design for the RIS phase shifts to maximize the number of users supported by HAPS-RIS while satisfying the rate requirements for each user in the HAPS-RIS zone. We practically relax and redesign the leader problem into an optimization problem with the lower computational complexity, and solve it with the proposed dynamic radius-based zone association with RIS clustering (DyRaZARC) technique. In the follower optimization problem, we aim to minimize the number of UAVs while meeting the rate requirement for each user in the UAV zone. To solve the follower problem, we propose a novel geometry-informed machine learning method, called $k$-means adaptive dynamic UAV selection (KADUS) technique. Our findings demonstrate that enhancing the number of RIS elements can significantly reduce the required number of UAVs, revealing a key strategy for optimizing aerial operations.
II Proposed Network Architecture
@@
We consider a downlink transmission scenario, shown in Fig. 1, where heterogeneous non-terrestrial infrastructures are exploited to bring connectivity to a remote area. The set of stationary users $\mathcal{I}$ is defined as $\mathcal{I}$={1, …, II}, where $i\in\mathcal{I}$ represents an individual user indexed by $i$. The users are uniformly distributed in a circular area where there is no nearby terrestrial infrastructure. The set of UAVs $\mathcal{J}$ is defined as $\mathcal{J}$ = {1, …, JJ}, where $j\in\mathcal{J}$ represents an individual UAV indexed by $j$. In this scenario, the UAVs remain stationary, hovering in place to provide service to the fixed users. The aim is minimizing the number of required UAVs with exploiting terrestrial CS and HAPS-RIS while providing full coverage with meeting the minimum rate requirements. In HAPS-RIS scenarios, CS uses high gain directional antenna, creating a dominated line of sight (LoS) connection with HAPS [8, 9, 10, 11, 12]. We consider UAVs and CS operate independently to provide connectivity and they are not connected to each other. All UAVs can be connected to the core network via a backhaul link. Due to heavy blockage, there is no direct link between the users and the CS; communication instead occurs through a cascaded channel through HAPS-RIS. The set of RIS elements $\mathcal{M}$ is defined as $\mathcal{M}$ = {1, …, MM}, where $m\in\mathcal{M}$ represents the RIS element indexed by $m$. Users can be connected to the network either through the UAV system or the HAPS-RIS setup, but not through both at the same time. The coverage area is a circle with the radius $R_{0}$ with the centre of $(x_{0},y_{0},z_{0})$, the CS location is $(x^{\rm{CS}},y^{\rm{CS}},z^{\rm{CS}})$, the location of the HAPS is $(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}})$, and the location of the UAV $j$ is $(x_{j}^{\rm{UAV}},y_{j}^{\rm{UAV}},z_{j}^{\rm{UAV}})$ in the Cartesian coordinate system. As can be seen in Fig. 1, there are two zones in the coverage area, i.e., the inner circle with the radius $R$ and the ring around it. The UAV zone is the area where the users are covered by UAVs and the HAPS-RIS zone is the ring area between the inner and outer circles, in which the users are covered by HAPS-RIS. The border which separates the UAV zone and the HAPS-RIS zone is called zone boundary. $\mathcal{B}$ is the set which includes the users in the UAV zone. Thus, $\mathcal{C}$$\mathcal{I-B}$ is the set which includes users in the HAPS-RIS zone. The set of equally spaced orthogonal subcarriers $\mathscr{L}$ is defined as $\mathscr{L}$={1, …, $L^{\rm{tot}}$}, where $l\in\mathscr{L}$ represents the subcarrier indexed by $l$. Let us consider the available bandwidth of $BW$ which is divided into $L^{\rm tot}$. Therefore, the bandwidth allocated to each orthogonal subcarrier is $B_{l}=\frac{BW}{L^{\rm{tot}}}$.
Figure 1: The Proposed Network Architecture
III Problem Formulation and Proposed Solution Techniques
@@
III-A Proposed HAPS-RIS Zone Design
@@
$\begin{array}[]{l}h^{\rm{CS}}_{im}=\\ \sqrt{G^{\rm{CS}}G_{i}\left(\mathcal{L}^{\rm{CS-RIS}}_{m}\right)^{-1}\left(\mathcal{L}^{\rm{RIS-user}}_{im}\right)^{-1}},\forall m,\forall i\in\mathcal{C},\end{array}$ (1)
where $G^{\rm{CS}}$ is the antenna gain of the CS and $G_{i}$ is the antenna gain of the user $i$. $(x^{\rm{CS}},y^{\rm{CS}},z^{\rm{CS}})$, $\left(x_{i},y_{i},z_{i}\right)$, $(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}})$, $\left(x^{\rm{RIS}}_{m},y^{\rm{RIS}}_{m},z^{\rm{RIS}}_{m}\right)$ are the CS, user $i$, HAPS, and RIS element $m$, coordinates, respectively. Using the Friis model, [13], the path loss between CS and RIS element $m$, $\forall m$, can be formulated as
$\begin{array}[]{l}\mathcal{L}^{\rm{CS},\rm{RIS}}_{m}=(\frac{4\pi f_{\rm c}}{c})^{2}\times\\ ((x^{\rm{RIS}}_{m}-x^{\rm{CS}})^{2}+(y^{\rm{RIS}}_{m}-y^{\rm{CS}})^{2}+(z^{\rm{RIS}}_{m}-z^{\rm{CS}})^{2}),\end{array}$ (2)
and the path loss between RIS element $m$ and the user $i$, $\forall m,\forall i\in\mathcal{C}$, can be formulated as
$\begin{array}[]{l}\mathcal{L}^{\rm{RIS},\rm{user}}_{im}=\left(\frac{4\pi f_{\rm{c}}}{c}\right)^{2}\times\\ ((x^{\rm{RIS}}_{m}-x_{i})^{2}+(y^{\rm{RIS}}_{m}-y_{i})^{2}+(z^{\rm{RIS}}_{m}-z_{i})^{2}),\end{array}$ (3)
@@
where $L^{\rm{CS}}$$\frac{L^{\rm{tot}}}{2}$ is available subcarriers, $\gamma^{l,\rm{CS}}_{i}$, the signal to noise ratio (SNR) of the received signal at user $i$ and subcarrier $l$, can be obtained as
@@
where $M$ is the total number of RIS elements, $N_{0}$ is the noise power spectral density, and $P^{l,\rm{CS}}$ is the CS transmit power allocated uniformly to the subcarrier $l$. $\rho^{l,\rm{CS}}_{im}$ is the binary variable which demonstrates the user $i$ is assigned to RIS element $m$ over subcarrier $l$.
Considering $\vartheta_{im}$ the reflection coefficient of the RIS element $m$ corresponding to the user $i$ as
@@
Remark 1: Users in the HAPS-RIS zone are assumed to be orthogonal to each other and to the users of the UAV zone. Hence, there is no inter-user interference.
Definition 1: Similar to the approach in [11], where the justification is elaborated in detail, we consider each user in the HAPS-RIS zone is allocated to its own group of subcarriers and RIS elements. The group of RIS elements which are allocated to a user is called RIS cluster. Given this assumption, $\rho^{l,\rm{CS}}_{im}$ can be obtained based on the value of $R$, the radius of the inner circle, which determines the zone boundary, elaborated in the proposed DyRaZARC Technique. Taking all these factors into account, the leader optimization problem can be formulated as
$\hskip-113.81102pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{OP 1}:\penalty 10000\ \penalty 10000\ \penalty 10000\ \max_{{R},\vartheta_{im}}\ U^{\rm{HAPS}}}$ (7a)
$\hskip-36.98866pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm{s}}.{\rm{t}}.:}\penalty 10000\ \penalty 10000\ r^{\rm{CS}}_{i}(R,\vartheta_{im})\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},}$ (7b)
@@
$\hskip-113.81102pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{OP 1-a}:\penalty 10000\ \penalty 10000\ \penalty 10000\ \max_{{R}}\ U^{\rm{HAPS}}}$ (8a)
$\hskip-36.98866pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm{s}}.{\rm{t}}.:}\penalty 10000\ \penalty 10000\ r^{\rm{CS}}_{i}(R)\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},}$ (8b)
where $\phi^{\ast}_{m}=\xi_{0}+\omega_{0}+2k\pi,\penalty 10000\ \forall m$ and $k$ can be an integer value.
@@
In the proposed geometrical case, the dimensions of RIS can be negligible compared to the altitude of HAPS, as it does not have much effect on the performance. Hence, without loss of generality, we relax the location of the RIS elements as
$\begin{matrix}\left(x^{\rm{RIS}}_{m},y^{\rm{RIS}}_{m},z^{\rm{RIS}}_{m}\right)\approx(x^{\rm{HAPS}},y^{\rm{HAPS}},z^{\rm{HAPS}}),&\forall m\end{matrix},$ (9)
the dimensions of the RIS and the coverage area is much less than the range of the distance between HAPS-RIS and the terrestrial components. [^1] The dimensions of the RIS are in the range of 100 meters while the distance between the HAPS-RIS and the terrestrial network is in the range of more than 20 kilometers. Accordingly, without loss of generality, we can relax the corresponding phases to fixed values as $\xi^{\rm{CS}}_{m}\approx\xi_{0},\forall m$ and $\omega_{im}\approx\omega_{0},\forall i\in\mathcal{C},\forall m,$ so the reflection gain of RIS element $m$ can be relaxed as $\vartheta_{im}\approx{\tilde{\vartheta}_{m}},\forall i\in\mathcal{C},\forall m$ where
@@
As (10) is independent from user index, $r^{\rm{CS}}_{i}(R,\tilde{\vartheta}_{m})$ can be relaxed to its upper bound by considering $\phi^{\ast}_{m}=$$\arg\max_{\phi_{m}}$${\tilde{\vartheta}_{m}}=\xi_{0}+\omega_{0}+2k\pi,\forall m$. Accordingly, (7b) is relaxed into (8b) and hence OP 1 is relaxed into OP 1-a. The proposed technique significantly reduces the computational complexity by eliminating the need to individually optimize the phase shifts for each RIS element. Given a massive number of RIS elements in HAPS-RIS scenarios, ranging from $10^{5}$ to $10^{8}$, see Section. IV, the proposed design can pave the way for scalability and enables practical implementation.
@@
$\hskip-113.81102pt\textbf{OP 1-b}:\penalty 10000\ \penalty 10000\ \penalty 10000\ \min_{R}\ R^{2}$ (11a)
$\hskip-28.45274pt{{\rm{s}}.{\rm{t}}.:}\penalty 10000\ \penalty 10000\ r^{\rm{CS}}_{i}(R)\geq r^{\rm{user}}_{0},\forall i\in\mathcal{C},$ (11b)